OpenStudy (anonymous):

Can someone help me on this question please, Simplify:(x-4sqroot(x) +3)/(x^2-1) Answer: (-3+sqrt(x))/((1+sqrt(x)) (1+x)) Thanks!

OpenStudy (anonymous):

\[Simplify: (x-4\sqrt{x}+3)/(x^2-1)\]

OpenStudy (nikvist):

\[\frac{x-4\sqrt{x}+3}{x^2-1}=\frac{(\sqrt{x}-1)(\sqrt{x}-3)}{x^2-1}=\frac{(\sqrt{x}-1)(\sqrt{x}-3)}{(\sqrt{x}-1)(\sqrt{x}+1)(x+1)}=\frac{\sqrt{x}-3}{(\sqrt{x}+1)(x+1)}\]